### v1

@@ 4 Formal Verification (FV) and Bounded Model Checking (BMC)
As manually labelling the entire dataset is unfeasible for such a large corpus of data, we employ a Formal Verification (FV) method, called Bounded Model Checking (BMC), to accurately detect vulnerabilities. In contrast to traditional static analysis tools, which often yield a high incidence of false positives due to reliance on pattern recognition without mathematical grounding [100], BMC offers rigorous validation. The Efficient SMT-based Bounded Model Checker (ESBMC) [14] effectively minimizes false positives and negatives. This approach ensures a fair comparison between programs produced by different LLMs.
4.1 State Transition System
@@
4.2 Bounded Model Checking
BMC is employed in FV to ascertain the correctness of a system up to a finite number of steps. This approach models the system as a finite state transition system and methodically examines its state space to a predefined depth. Recent BMC modules are capable of processing a variety of programming languages such as C, C++, JAVA, or Kotlin [101, 100, 102, 103, 104, 105, 106]. The process starts with the program code from which a CFG is derived. In this CFG, nodes represent deterministic or nondeterministic assignments or conditional statements, while edges indicate potential changes in the program’s control flow. Essentially, each node is a block that encapsulates a set of instructions with a unique entry and exit point, and edges indicate potential transitions to other blocks. The CFG is then converted into Static Single Assignment (SSA) form and further into a State Transition System (STS), which a Satisfiability Modulo Theories (SMT) solver can interpret. The SMT solver checks if a given formula, representing the program’s correctness within a bound $k$, is satisfiable, indicating the existence of a counterexample to the properties being verified. If no errors are found and the formula is unsatisfiable within the bound $k$, it suggests the program has no vulnerabilities within that limit. Thus, if the solver concludes satisfiability within a bound $\leq k$, it confirms the presence of a vulnerability through a counterexample.
@@ 4.2.1 The Bounded Model Checking Problem
Herein, $I$ denotes the initial state set of $\mathcal{ST}$, and $T(s_{i},s_{i+1})$ embodies the transition relation within $\mathcal{ST}$ between consecutive time steps $i$ and $i+1$. Thus, the logical expression $I(s_{0})\wedge\bigwedge^{k-1}_{i=1}T(s_{i},s_{i+1})$ depicts the execution pathways of $\mathcal{ST}$ spanning a length $k$, and $BMC_{\Phi}(k)$ can be satisfied if and only if for some $i\leq k$ there exists a reachable state at time step $i$ in which $\phi$ is violated. If $BMC_{\Phi}(k)$ is satisfied, it implies a violation of $\phi$, permitting an SMT solver to deduce a satisfying assignment from which the program variables’ values can be derived to assemble a counterexample. By definition, a counterexample, or trace, for a violated property $\phi$, is defined as a finite sequence of states $s_{0},\ldots,s_{k}$, where $s_{0},\ldots,s_{k}\in S$ and $T(s_{i},s_{i+1})$ holds for $0\leq i<k$. These counterexamples hold significant importance for us, as we explicitly seek out these violations to compare and determine which code generated by LLMs is “more secure”. Fewer violated properties indicate that the LLM can produce more secure code. In this context, it is important to note the influence of cyclomatic complexity. Merely experiencing fewer errors with an LLM does not necessarily indicate superiority; it could simply be generating less intricate problems. Therefore, assessing both property violations and cyclomatic complexity concurrently offers a more comprehensive understanding of an LLM’s proficiency in secure coding.
If the Equation (1) is unsatisfiable, it implicates no error state as reachable within $k$ steps or fewer, hence no software vulnerability exists within the bound $k$. By searching for counterexamples within this bound, we can establish, based on mathematical proofs, whether a counterexample exists and whether our program $\mathcal{P}$ contains a security vulnerability. This method detects security concerns such as buffer overflows and division by zero or null dereference failures. It is worth noting that if a program is classified as vulnerable, this assessment relies on counterexamples, effectively eliminating the possibility of false positives. Conversely, in situations where no counterexample exists, we can confidently assert that the program is free from vulnerabilities up to the bound $k$, ensuring the absence of false negatives. By adopting this strategy, we aim to classify each program by detecting violated properties in the generated code. While this method offers greater precision and improved outcomes compared to standard static analysis tools, it is also more time-consuming, demands significant resources, and is limited to identifying only a predefined set of vulnerabilities.
@@ 4.3 Efficient SMT-based Context-Bounded Model Checker
Annually, the International Competition on Software Verification, known as SV-COMP, challenges various programs to detect bugs and ensure software safety. ESBMC excels in this competition by solving the highest number of verification tasks within a strict $10$-$30$ second time limit per program, as evidenced in SV-COMP 2023 [107]. Given its performance, ESBMC was selected as our primary BMC tool. As a robust, open-source model checker for C/C++, Kotlin, and Solidity programs, ESBMC addresses a wide range of safety properties and program assertions, including out-of-bounds array access, illegal pointer dereference, integer overflows, and memory leaks. It employs advanced verification techniques such as incremental BMC and k-induction, supported by state-of-the-art SMT and Constraint Programming (CP) solvers. ESBMC’s effectiveness in falsification and bug-finding is underscored by its multiple accolades in SV-COMP, including 6 gold, 4 silver, and 10 bronze medals.
4.3.1 Unbounded model checking using ESBMC
In the original work of [9], the bounded model checking (BMC) problem was addressed with a bound set to $k=1$. Although we have observed that most vulnerabilities can be detected with these settings empirically, the nature of BMC can sometimes lead to erroneous conclusions. For example, if a property violation occurs at level $k=2$, then $BMC_{\Phi}(1)$ will fail to detect the bug and report the verification as successful, giving a false impression. As a result, in the FormAI-v1 dataset, numerous samples were previously classified as ”NON-VULNERABLE up to bound $k$. We have transitioned from bounded to unbounded model checking to capture more vulnerabilities or prove their absence for each sample. This means that the program is incrementally unwound until a bug is found or until the completeness threshold is reached, i.e., states corresponding to the program terminating. This incremental BMC approach ensures that smaller problems are solved sequentially instead of guessing an upper bound for the verification. Applying these settings, we have successfully identified more vulnerabilities or prove that no vulnerability exists. Consequently, if the verification process is completed successfully, we can conclude that the program has no violated properties. While this approach requires significantly more computational power, it has proven effective in revealing a greater number of vulnerabilities or proving their absence, as we will demonstrate in Section 6.
@@ 5 Methodology and Dataset Creation
The FormAI-v2 dataset consists of two main components: the AI-generated C programs and their vulnerability labeling. In the data generation phase, we have created $265,000$ samples. This paper implements a similar methodology introduced in the original work [9]. The main difference is that we employ eight different LLMs rather than solely GPT-3.5. Moreover, as discussed in the previous section, we have adopted a more expansive unbounded verification model. This section covers the dataset generation methodology and the specific prompts for producing the C code samples. The overview of the generation and vulnerability labeling mechanism is depicted in Figure 5.
@@ 5.1 Code Generation
During the creation process, special attention was given to ensure the diversity of the FormAI-v2 dataset, which contains $265,000$ compilable C samples. Using a basic prompt like ”generate a C program” often results in similar outputs, such as programs that add two numbers or perform simple string manipulations, which do not meet our objectives. We need to systematically generate a comprehensive and varied collection of small programs. To meet this goal, we have developed a prompting method consisting a dynamic and a static part. The static component remains consistent and unchanged, while the dynamic portion of the prompt undergoes continuous variation. An example of how a single prompt looks is shown under Figure 6.
Figure 6: Dynamic code generation prompt
The dynamic part of the prompt, highlighted as [Type] and [Style], represent distinct categories within the prompt, each encompassing different elements. Every API call randomly selects a Type category from a set of $200$ elements. This category contains topics such as Wi-Fi Signal Strength Analyzer, QR code reader, Image Steganography, Pixel Art Generator, Scientific Calculator Implementation, etc. Similarly, a coding Style is chosen from a set of $100$ elements during each query. This helps minimize repetition, as coding styles such as “excited”, “relaxed”, or “mathematical” are randomly combined with a Type category. By employing this method, we can generate $200\times 100=20,000$ distinct combinations. As demonstrated by insights from [108, 80], there’s a need for a code base that supports diverse settings while ensuring tasks remain concise to fit within the token constraints of large language models (LLMs). With this consideration, we designed tasks in the Type category, selecting prompts that LLMs can efficiently process. For instance, complex prompts like “Create a CRUD application using React for the front-end, Node.js with Express for the back-end, and MongoDB for the database” must be broken down into smaller, manageable tasks. Furthermore, tasks with different styles, such as ”File handling” with a ”thriller” versus a ”happy” style, lead to different representations in the vector space upon tokenization. Despite potential compatibility issues between certain Type-Style combinations, encouraging LLMs to code in varied styles has generally enhanced the diversity and distinctiveness of responses to identical tasks.
@@
Once a C code is generated, the GNU C compiler [^5] https://gcc.gnu.org is employed to verify whether the corresponding code is compilable. During the code generation process, we ensure that the FormAI-v2 dataset exclusively consists of compilable code while excluding any other code that does not meet this criterion. Different models can generate varying percentages of compilable code depending on their parameter size. Models like GPT-4, GEMINi-pro, or FALCON-180B can achieve compilation rates higher than 90%, whereas smaller models with 7B parameters typically produce C code with a compilability rate between 55-70%. The primary reason for having non-compilable code was due to the absence of necessary headers, such as math.h, ctype.h, or stdlib.h. The generation of $265,000$ code samples was completed within $48$ hours. As the cost of generation associated with different models can significantly vary, we did not generate the same number of samples from each model. For example, GPT-4 API calls can be up to $60$ times as expensive as GPT-3.5. Table 2 shows how many samples we acquired from each LLM.
Table 2: Content of the FormAI v2.0 dataset.
@@
GPT-4 OpenAI N/A Proprietary 1000
Llama2-13B Meta 13B Open 5000
Mistral-7B Mistral AI 7B Apache 2.0 10000
CodeLlama-13B Meta 13B Proprietary 12000
Gemini 1.0 Pro Google N/A Proprietary 40000
Gemma-7b Google 7B Gemma-terms-of-use 47000
Falcon-180B TII 180B Apache 2.0 72000
GPT-3.5 OpenAI 175B Proprietary 78000

### v2

@@ 4 Formal Verification (FV) and Bounded Model Checking (BMC)
Before presenting the methodology used to construct the dataset and examining the performance of different LLMs, this section will introduce key Formal Verification (FV) concepts to clarify the approach adopted in developing the dataset. Since manually labeling the entire dataset is not feasible for such a large volume of data, we use an FV technique known as Bounded Model Checking (BMC) to detect vulnerabilities in the generated C samples precisely. In contrast to traditional static analysis tools, which frequently produce a high number of false positives due to their reliance on pattern recognition without a solid mathematical foundation [105], BMC provides rigorous validation that can help minimize both false positives and false negatives in the findings.
4.1 Preliminaries for the Data Labeling Method
To enhance understanding and ensure the reproducibility of our methodology, we introduce some key definitions, including State Transition Systems (STS), the BMC problem, and the specific tools chosen for our labeling method, considering the many FV tools available in the market.
4.1.1 State Transition System
@@
4.1.2 Bounded Model Checking
BMC is employed in FV to ascertain the correctness of a system up to a finite number of steps. This approach models the system as a finite state transition system and methodically examines its state space to a predefined depth. Recent BMC modules are capable of processing a variety of programming languages such as C, C++, JAVA, or Kotlin [106, 105, 107, 108, 109, 110, 111]. The process begins with the program code, from which a control-flow graph (CFG) [112] is derived. In this CFG, nodes represent deterministic or nondeterministic assignments or conditional statements, while edges indicate potential changes in the program’s control flow.
Essentially, each node is a block that encapsulates a set of instructions with a unique entry and exit point, and edges indicate potential transitions to other blocks. The CFG is then converted into Static Single Assignment (SSA) form and further into a State Transition System (STS), which a Satisfiability Modulo Theories (SMT) solver can interpret. The SMT solver checks if a given formula, representing the program’s correctness within a bound $k$, is satisfiable, indicating the existence of a potential counterexample to the properties being verified. If no errors are found and the formula is unsatisfiable within the bound $k$, it suggests the program has no vulnerabilities within that limit. Thus, if the solver concludes satisfiability within a bound $\leq k$, it confirms the presence of a vulnerability through a counterexample.
@@ 4.1.3 The Bounded Model Checking Problem
Herein, $I$ denotes the initial state set of $\mathcal{ST}$, and $T(s_{i},s_{i+1})$ embodies the transition relation within $\mathcal{ST}$ between consecutive time steps $i$ and $i+1$. Thus, the logical expression $I(s_{0})\wedge\bigwedge^{k-1}_{i=1}T(s_{i},s_{i+1})$ depicts the execution pathways of $\mathcal{ST}$ spanning a length $k$, and $BMC_{\Phi}(k)$ can be satisfied if and only if for some $i\leq k$ there exists a reachable state at time step $i$ in which $\phi$ is violated. If $BMC_{\Phi}(k)$ is satisfied, it implies a violation of $\phi$, permitting an SMT solver to deduce a satisfying assignment from which the program variables’ values can be derived to assemble a counterexample. By definition, a counterexample, or trace, for a violated property $\phi$, is defined as a finite sequence of states $s_{0},\ldots,s_{k}$, where $s_{0},\ldots,s_{k}\in S$ and $T(s_{i},s_{i+1})$ holds for $0\leq i<k$. These counterexamples hold significant importance for us, as we explicitly seek out these violations to compare and determine which code generated by LLMs is “more secure”. Fewer violated properties indicate that the LLM can produce more secure code.
In this context, it’s important to note that fewer errors in a C program generated by an LLM do not necessarily indicate superiority; the model may simply be producing simpler, shorter programs. Therefore, evaluating both property violations and code complexity metrics, such as Source Lines of Code (SLOC) or Cyclomatic Complexity (CC)  [17], can be a good starting point to determine the complexity of the generated programs. For example, a basic “print hello world” program will not contain any vulnerabilities, but that doesn’t mean it’s a good program. This is why metrics like SLOC (Source Lines of Code) and CC (Cyclomatic Complexity) are crucial, as they help identify overly simplistic or short code that may have fewer vulnerabilities simply due to its simplicity, not because it’s well-written.
If the Equation (1) is unsatisfiable, it implicates no error state as reachable within $k$ steps or fewer. Hence, no software vulnerability exists within the bound $k$. By searching for counterexamples within this bound, we can establish, based on mathematical proofs, whether a counterexample exists and whether our program $\mathcal{P}$ contains a security vulnerability. This method detects security issues such as buffer overflows, division by zero, and null dereference failures. Notably, if a program is identified as vulnerable, this determination is based on counterexamples, effectively reducing the likelihood of false positives. Conversely, in cases where no counterexample is found, we can confidently state that the program is free from vulnerabilities up to the bound $k$, thereby minimizing false negatives. By adopting this strategy, we aim to classify each program by detecting violated properties in the generated code.
@@ 4.2 Efficient SMT-based Context-Bounded Model Checker
Numerous BMC tools could meet our needs. However, we aimed to select a tool offering high performance and detection rates. Annually, the International Competition on Software Verification, known as SV-COMP, challenges various programs to detect bugs and ensure software safety. In this competition, the Efficient SMT-based Bounded Model Checker (ESBMC) [14] stands out by solving the highest number of verification tasks within a strict 10-30 second time limit per program, as demonstrated in SV-COMP 2023 [113].
@@
Given its performance, ESBMC was selected as our primary BMC tool. As a robust, open-source model checker for C/C++, Kotlin, and Solidity programs, ESBMC addresses a wide range of safety properties and program assertions, including out-of-bounds array access, illegal pointer dereference, integer overflows, and memory leaks. It employs advanced verification techniques such as incremental BMC and k-induction, supported by state-of-the-art SMT and Constraint Programming (CP) solvers. ESBMC’s effectiveness in bug-finding is highlighted by its numerous achievements in SV-COMP, earning 6 gold, 4 silver, and 10 bronze medals.
4.2.1 Identifiable Bugs Using ESBMC
Although using ESBMC to identify bugs provides greater precision than traditional static analysis tools, it is also more time-consuming, requires substantial resources, and is limited to detecting a specific set of vulnerabilities. This raises a natural question: what types of vulnerabilities can BMC detect, and which are the ones it cannot?
BMCs primarily address low-level, code-centric issues such as buffer overflows, memory leaks, and assertion failures. For instance, ESBMC identifies software errors by simulating a limited portion of a program’s execution with all possible inputs. However, vulnerabilities such as SQL injection, code injection, and XSS generally fall outside the scope of BMCs because creating a general mathematical model to represent how a web browser or database interprets code is highly challenging. Additionally, SQL queries and HTML scripts can be written in various ways, making it impossible to create an exact abstract formula. This is particularly problematic because the primary goal of formal verification is to model all possible inputs to verify the system and identify property violations effectively.
When using ESBMC, the verification result of each C program falls into one of four major categories: Verification Success ($\mathcal{VS}$), Verification Failed ($\mathcal{VF}$), Verification Unknown ($\mathcal{VU}$), and Parsing Errors ($\mathcal{ER}$), as illustrated in Table 2. These categories are mutually exclusive, meaning a single sample cannot belong to more than one category.
Table 2: The Four Main Categories for Vulnerability Classification With ESBMC
Category Description
$\mathcal{VS}$: Verification Success The set of samples for which the verification process was completed successfully with no vulnerabilities detected.
$\mathcal{VU}$: Verification Unknown        (Timeout) The set of samples for which the verification process did not complete within the allotted time frame. Although no counterexample was found within the time limit, this does not guarantee the absence of vulnerabilities in the program with a longer time frame; therefore, the verification status remains unknown.
$\mathcal{ER}$: Error The set of samples for which the verification status resulted in an error. This typically occurs due to a parsing error in ESBMC, an issue in the GOTO converter, or other problems with the SMT solver.
The most relevant category for our analysis is Verification Failed ($\mathcal{VF}$), which can be further subdivided into five main types: dereference failures ($\mathcal{DF}$), arithmetic overflows ($\mathcal{AO}$), buffer overflows ($\mathcal{BO}$), array bounds violations ($\mathcal{ABV}$), and other miscellaneous vulnerabilities ($\mathcal{MV}$). These five categories encompass 33 subcategories that ESBMC can identify, as illustrated in Table LABEL:tab:vulnerability.
It is important to note that when we refer to “Verification Successful” ($\mathcal{VS}$), it indicates that the specific bugs listed in Table LABEL:tab:vulnerability were not detected in the programs. However, this does not rule out the presence of other types of vulnerabilities in these programs, such as command injection, cryptographic weaknesses, SQL injection, and similar issues. Table LABEL:tab:vulnerability also displays each vulnerability’s corresponding Common Weakness Enumeration (CWE) number. It is important to note that ESBMC does not provide an exact mapping of vulnerabilities to CWE numbers; the mapping presented here was performed manually.
The multifaceted nature of software flaws often results in a single vulnerability associated with multiple CWE identifiers. Table LABEL:tab:vulnerability categorizes the most common vulnerabilities and the corresponding CWEs identified within these categories. In total, $42$ unique CWE were identified in the dataset. From MITRE’s Top 25 Most Dangerous Software Weaknesses for 2023 list, six is present in our list as shown in Table 3.
Table 3: CWEs From 2023’s MITRE Top 25.
Rank CWE Description
1 CWE-787: Out-of-bounds Write
4 CWE-416: Use After Free
6 CWE-20: Improper Input Validation
7 CWE-125: Out-of-bounds Read
12 CWE-476: NULL Pointer Dereference
14 CWE-190: Integer Overflow or Wraparound
The remaining CWEs in the top 25 list are related to web vulnerabilities like SQL injection, XSS, and authentication, which are irrelevant to our C language samples. It is vital to emphasize that, in our situation, classifying the C programs based on CWE identifiers is not practical, contrary to what is done for other databases like Juliet. As shown in Table 1, most datasets contain only one vulnerability per sample. In the datasets ReVeal, BigVul, and Diversevul, a function is vulnerable if the vulnerability-fixing commit changes it, while in Juliet, a single vulnerability is introduced for each program.
In FormAI, a single file often contains multiple vulnerabilities. As noted, a single vulnerability can be associated with multiple CWEs. Additionally, multiple CWEs can be required for a vulnerability to be exploitable. As an example, “CWE-120: Buffer Copy without Checking Size of Input (Classic Buffer Overflow)”, can happen as a result of “CWE-676: Use of Potentially Dangerous Function”, which can be the scanf function. If this is combined with “CWE-20: Improper Input Validation”, it can result in “CWE-787: Out-of-bounds Write”. Labeling the vulnerable function name, line number, and vulnerability type identified by the ESBMC module provides granular information that can benefit machine learning training. This level of detail can allow models to discern patterns and correlations with higher precision, thereby improving vulnerability prediction and detection capabilities.
Table 4: Detailed Categorization of Vulnerabilities Detected by ESBMC
Description CWE Associated CWEs
$\mathcal{DF}$: Dereference failures:
 1. NULL pointer CWE-476 CWE-690, CWE-391
 2. Invalid pointer CWE-822 CWE-119, CWE-787, CWE-822
 3. Forgotten memory CWE-825 CWE-401, CWE-404, CWE-459
 4. Array bounds violated CWE-125 CWE-119, CWE-787
 5. Invalidated dynamic object CWE-824 CWE-416, CWE-415
 6. Access to object out of bounds CWE-125 CWE-119, CWE-787
 7. Accessed expired variable pointer CWE-416 CWE-825
 8. Write access to string constant CWE-843 CWE-758
 9. Of non-dynamic memory CWE-590 CWE-415, CWE-415, CWE-762
 10. IBTA CWE-843 CWE-119
 11. Oversized field offset CWE-787 CWE-119, CWE-125, CWE-823
 12. Data object accessed with code type CWE-843 CWE-686, CWE-704
$\mathcal{AO}$: Arithmetic overflows:
 13. On sub CWE-191 CWE-20, CWE-190, CWE-192
 14. On add CWE-190 CWE-20, CWE-191, CWE-192
 15. On mul CWE-190 CWE-20, CWE-191, CWE-192
 16. Floating-point ieee_mul CWE-190 CWE-681
 17. Floating-point ieee_div CWE-682 CWE-369, CWE-681
 18. Floating-point ieee_add CWE-190 CWE-681
 19. Floating-point ieee_sub CWE-190 CWE-681
 20. On div CWE-190 CWE-20, CWE-369
 21. On shl CWE-190 CWE-192
 22. On modulus CWE-190 CWE-20, CWE-191
 23. On neg CWE-191 CWE-190, CWE-192
$\mathcal{BO}$: Buffer overflow:
 24. On scanf CWE-120 {CWE-20, CWE-121, CWE-122
 25. On fscanf CWE-120 CWE-129, CWE-131, CWE-628
 26. On sscanf CWE-120 CWE-676, CWE-680, CWE-787}
$\mathcal{ABV}$: Array bounds violations:
 27. lower bound CWE-129 {CWE-119, CWE-125, CWE-129
 28. upper bound CWE-788 CWE-131, CWE-193, CWE-787}
 29. VLA array size in bytes overflows CWE-190 CWE-131, CWE-680
$\mathcal{MV}$: Miscellaneous Vulnerabilities:
 30. Division by zero CWE-369 CWE-691
 31. The pointer to a file must be valid CWE-476 CWE-690, CWE-459
 32. Same object violation CWE-628 CWE-843, CWE-668
 33. ZOFO CWE-761 CWE-415, CWE-590
Legend:
ZOFO: Operand of free must have zero pointer offset, IBTA: Object accessed with incompatible base type
Since our programs exhibit numerous vulnerabilities, including multiple occurrences of the same type, categorizing each solely into one CWE group, as seen with Juliet, would be sub-optimal for training purposes. This method fails to communicate crucial details about the vulnerabilities. For instance, both “Arithmetic overflow on add” and “Arithmetic overflow on div” are assigned the same primary CWE, manifesting differently in the source code. Therefore, merely labeling them with CWEs does not offer the same level of granularity and makes the dataset less suitable for ML.
While other datasets focus more on CWEs related to vulnerabilities that could be exploited, ESBMC also detects issues related to software safety. For this reason, in the FormAI dataset, we did not assign a single CWE to each vulnerability. However, based on our mapping in Table LABEL:tab:vulnerability, one can easily associate an ESBMC vulnerability with the closest CWE number if needed.
@@ 5 Methodology and Dataset Creation
Figure 4 provides an overview of the generation and vulnerability labeling mechanism for the FormAI-v2 dataset. This process is divided into two main components: the C program generation (consisting of 1. C program generation using different LLMs and 2. dataset preprocessing) and the classification (including 3. ESBMC classification and 4. dataset creation).
@@ 5.1 Code Generation
During the creation process, special attention was given to ensure the diversity of the FormAI-v2 dataset, which contains 331 000331000331\,000331 000 compilable C samples. Using a prompt like “generate a C program” repeatedly, would yields similar outputs, such as adding two numbers or simple string manipulations, which does not satisfy our objectives. Instead, our goal is to generate a diverse and comprehensive set of small programs. To meet this, we have developed systematic prompting method consisting a dynamic and a static part. The static component remains unchanged for all prompts, while the dynamic portion undergoes continuous variation. An example of how our prompt template looks like is shown under Figure 5.
Figure 5: Dynamic Code Generation Prompt.
The dynamic part of the prompt, highlighted as [Type] and [Style], represent distinct categories within the prompt, each encompassing different elements. Every API call randomly selects a Type category from a set of $200$ elements. This category contains topics such as Wi-Fi Signal Strength Analyzer, QR code reader, Image Steganography, Pixel Art Generator, Scientific Calculator Implementation, etc. Similarly, a coding Style is chosen from a set of $100$ elements during each query. This helps minimize repetition, as coding styles such as “excited”, “relaxed”, or “mathematical” are randomly combined with a Type category. Our primary objective was to identify and capture as many vulnerabilities as possible. This method can generate 200×100=20 00020010020000200$\times$ 100=20\,$0$ 100 = 20 000 distinct combinations. As demonstrated by insights from [114, 86], there’s a need for a code base that supports diverse settings while ensuring tasks remain concise to fit within the token constraints of large language models (LLMs).
This raises a key question: If we generate a dataset of over 300,000 instances but only 20,000 distinct combinations, will it lead to redundancy? Will the same or different models produce identical outputs for these repeated prompts? To address this, we will conduct clone code detection in the next section to ensure the generated code is unique. Selecting prompts that LLMs can efficiently process is important, therefore we designed tasks in the Type category accordingly. For instance, complex prompts like “Create a CRUD application using React for the front-end, Node.js with Express for the back-end, and MongoDB for the database” must be broken down into smaller, manageable tasks. Furthermore, tasks with different styles, such as ’File handling’ with a ’romantic’ versus a ’happy’ style, lead to distinct outputs, which are reflected in different representations in the vector space upon tokenization. Despite potential compatibility issues between certain Type-Style combinations, encouraging LLMs to code in varied styles has generally enhanced the diversity of responses to identical Types.
@@
Once a C code is generated, the GNU C compiler [^5] https://gcc.gnu.org is employed to verify whether the corresponding code is compilable. During the code generation process, we ensure that the FormAI-v2 dataset exclusively consists of compilable code while excluding any other code that does not meet this criterion. Different models can generate varying percentages of compilable code depending on their parameter size. Models like GPT-4o-mini, Gemini Pro, or Falcon-180B can achieve compilation rates higher than 90%, whereas smaller models with 7B parameters typically produce C code with a compilability rate between 55-70%.
The primary reason for having non-compilable code was due to the absence of necessary headers, such as math.h, ctype.h, or stdlib.h. As the cost of generation associated with different models can significantly vary, we did not generate the same number of samples from each model. While some tested models are open source, their operational costs and GPU usage remain significant. For instance, running Falcon-180B on AWS can cost around 40 USD per hour. Table 5 presents the samples obtained from each LLM.
Table 5: Content of the FormAI-v2 Dataset.
@@
GPT-4o-mini OpenAI N/A Proprietary 40 0004000040\,00040 000
Llama2 13B Meta 13B Open 20 0002000020\,00020 000
Mistral-7B Mistral AI 7B Apache 2.0 10 0001000010\,00010 000
Code Llama 13B Meta 13B Proprietary 12 0001200012\,00012 000
Gemini Pro 1.0 Google N/A Proprietary 40 0004000040\,00040 000
Gemma-7B Google 7B Gemma-TOU 47 0004700047\,00047 000
Falcon-180B TII 180B Apache 2.0 72 0007200072\,00072 000
Falcon2-11B TII 11B Apache 2.0 12 0001200012\,00012 000
GPT-3.5-turbo OpenAI 175B Proprietary 78 0007800078\,00078 000
5.2 Clone Code Detection
For our purposes, we do not require entirely different programs, even if they are intended to accomplish the same task. Our goal is to observe how frequently models introduce vulnerabilities into the programs. Therefore, even small changes can be interesting, or if a model repeatedly makes the same coding errors which lead to different vulnerabilities. Additionally, minor variations help to generalize the dataset during machine learning training, so completely removing similar code is not our objective.
To measure how diverse the dataset is, we used a code clone detection mechanisms to flag highly similar files. Using the state-of-the-art tool NiCad 7 (Automated Detection of Near-Miss Intentional Clones)[18], we performed clone detection within individual tasks and across datasets generated by different models. NiCad can detect four distinct types of clones, each with varying thresholds. Type 1 clones are exact duplicates of code fragments, with no changes except for whitespace and comments. Type 2 clones permit minor modifications, such as renaming variables or altering formatting, while maintaining the original logic and structure.
Type 3-1 clones introduce greater flexibility, allowing small additions, deletions, or modifications while preserving overall functionality. Type 3-2 clones allow for even more substantial changes, but the core behavior of the code remains consistent. In our clone detection process for Type 3-1 and Type 3-2 clones, we applied a 10% threshold to eliminate similar codes. A larger threshold could eliminate valuable code fragments for machine learning, where small deviations may still result in the same vulnerability. Even slight variations can improve the training process by offering diverse representations of the same vulnerability, enabling the model to recognize it more effectively across various scenarios. Table 6 shows each dataset’s number of clones identified.
Table 6: Different Types of Clones Removed From the Dataset
LLM Model Sample size Type1 Type2 Type 3-1 Type 3-2 $\Delta$(%)
Falcon2-11B 12 0001200012\,00012 000 00 00 $1$ $36$ $0.30$
Mistral-7B 10 0001000010\,00010 000 $1$ $1$ $11$ $59$ $0.59$
CodeLlama-13B 12 0001200012\,00012 000 $3$ $5$ $12$ $128$ $1.10$
GPT-3.5-turbo 78 0007800078\,00078 000 $118$ $301$ $502$ 1 75617561\,7561 756 $2.25$
GPT-4o-mini 40 $0 00$ $24$ $31$ 1 07510751\,0751 075 $2.69$
Gemini Pro 1.0 40 0004000040\,00040 000 $12$ $150$ $871$ 25512551\,2551 255 $3.14$
Falcon-180B 72 0007200072\,00072 000 $42$ $363$ $541$ 3 46434643\,4643 464 $4.81$
Llama2-13B 20 0002000020\,00020 000 $165$ $607$ 1 00110011\,0011 001 2 21422142\,2142 214 $11.07$
Gemma-7B 47 0004700047\,00047 000 $657$ 3 22932293\,2293 229 2 99729972\,9972 997 10 1991019910\,19910 199 $21.70$
The last column, $\Delta(\%)$, shows the percentage of Type 3-2 clones that were detected and removed from the dataset. A higher percentage indicates that the LLM generated more similar, redundant code samples.
[…]
However, Type 2 is not a subset of Type 3-1 because the threshold for Type 2 clones is exactly zero—meaning only variable changes are allowed across the entire code with no additional modifications. In contrast, Type 3-1 allows for up to a $10\%$ modification threshold, which can include variable changes, deletions, additions, or structural modifications, as long as they remain within the $10\%$ limit.
The most flexible clone category is Type 3-2, where a $10\%$ threshold applies to the entire code without restrictions. This means that any kind of modification, including variable changes throughout the entire program, is allowed. To ensure the dataset’s quality, we removed all clones up to and including Type 3-2.
After filtering out these clones from each LLM-generated subset, we applied Type 3-2 detection to the entire dataset to identify any similar code across different models. This process revealed an additional $283$ Type 3-2 clones, which were subsequently removed. In total, 20 4692046920\,46920 469 programs were excluded, resulting in a final dataset of 310 531310531310\,531310 531 unique files. This demonstrates that the dataset is diverse, with only $6.18\%$ of the original programs being similar.